Status Report

ISS On-Orbit Status 13 Aug 2002

By SpaceRef Editor
August 13, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

After crew wake-up last night at 8:00 pm EDT (sleep time: 11:30 am), followed by the regular schedule of morning inspection, breakfast, work preparation and DPC (daily planning conference) with the ground, the pace of EVA preparations on board picked up rapidly.

First, all crewmembers completed a one-hour review of the detailed timeline for EVA-7.  [The “cyclogram” is a breakdown of the spacewalk in individual activities along the prescribed crew translation path out of the DC-1 airlock to the FGB and PMA-1 (to pick up the DPP debris shields bundle stored there) and thence back to the SM for the installation of the six panels around a portion of the conical hull section.  Transfer of the DPP bundle will be handled by means of the Strela-2  cargo crane (GSt-2) mounted on the DC-1.  The prescribed translation path is mapped in terms of numbered handrail locations, and special charts showing the numerous handrails on the FGB were uplinked for the timeline review.  Other objectives of the spacewalk are retrieval/replacement of the Kroma-1 research tablet (with collected thruster effluents) and a visual and photographic inspection of external SM surface locations.  EVA-7 (i.e., the 7th from the Pirs DC-1) will begin at 3:40 am EDT, 8/16, and is estimated to last 5h 55 min.]

Later, CDR Valery Korzun and FE-2 Sergey Treschev completed pressure checks of the SM BK-3 oxygen (O2) tanks and the BNP portable repress O2 tank in the DC-1 airlock module.

Treschev set up the pre-EVA communications and tracking configuration for the subsequent tests of Orlan systems.  [Orlan-M suit communications to the SM have two links: by hard-wire via umbilical from the DC-1 airlock, and by VHF radio (“Korona”).  During EVA, comm loops exist between SM and TsUP (MCC-Moscow) via NIP (Russian ground station) on VHF, as well as between the US segment and MCC-Houston via TDRS satellite and White Sands on S-band.  Both Mission Control Centers are of course interconnected.]

Both EV crewmembers checked out their medical electrocardiogram harnesses (PKO-BETA08) on the GAMMA-1 medical examination panel (PKO) before they set up for a thorough test of the Orlan telemetry and communications from the DC-1.  Other preparations involved leak checks and valve function tests on the Orlans and their support systems (BSS), as well as actuation verification of pressure equalization valves (KBD) from the EVA support panels (POV) in the PkhO and DC-1 modules.

Later in the day, Korzun and Whitson installed external EVA equipment on their Orlan suits and tagged up with TsUP EVA specialists via S-band.

Peggy Whitson reconnected the UOP (utility operations panel) bypass cable to the Lab RWS DCP (robotic workstation/display & control panel), and later MCC-H powered up the MSS (mobile service system), to perform some Robotics checkouts, viz., of the MCAS (mobile base system common attach system) latch, and the MBS (mobile base system) mast camera. [The MCAS activities were to verify latch mechanism functionality on both strings, and all commanding was performed from the ground, but Peggy needed to throw a safing switch as part of the evening DPC to allow the ground to switch strings.  The mast camera OCR (on-orbit checkout requirement) was recorded, and the tape will be returned to the ground.]

FE-1 Whitson performed the scheduled troubleshooting of the communications problem between the Airlock BSA (battery stowage assembly) with its battery chargers and the SSC (station support computer) reported before.  [In the morning she checked the SSC comm port (software and hardware), inspected the connecting comm cable visually, and installed/removed EVA batteries in BSA.  In the afternoon, she powered up BC3 (battery charger #3) and tested communications/discharge.]

Daily routine servicing tasks were performed by Whitson (autonomous Lab payload status check) and Treschev (BRPK-2 water condensate separator inspection, IMS delta file preparation).  

Peggy performed the weekly TVIS (treadmill) maintenance.  

Later, she also completed her share of the weekly NTXN “Interactions” data collection for the encrypted mood, group and journal questionnaire on the HRF (human research facility) PC.

All crewmembers performed their daily physical exercise on TVIS, RED and VELO with load trainer.

POC (Payload Operations Center/Huntsville) thanked Peggy for getting the SAMS (space acceleration measurement system) payload back in operation yesterday with a reboot of its ICU (interim control unit) in ER4 (EXPRESS rack #4).

During removal of SUBSA-07, the fifth processed sample of the SUBSA (solidification using a baffle in sealed ampoules) experiment from the furnace in the MSG (Microgravity Science Glovebox) on 8/11, the ampoule broke apart, still contained within the MSG.  Peggy did some minor clean-up, but shards and contents remain in the MSG, and the glovebox is out of service until POC has devised a suitable recovery scheme.  Due to the EVA activities, no research ops were scheduled during this and the next week for MSG (which clearly served its intended purpose well).

Both BGAs (beta gimbal assemblies) of the P6 photovoltaic power module continue in directed position (Blind) mode for the on-going dual-rate angle test.

Today’s targets for the US CEO (Crew Earth Observations) program were Borneo Fires (fires in western Borneo [left of track all the way to the limb] are making news as “brown haze” afflicts Indonesia and its neighbors. Hundreds of fires set to clear forest are the main culprits), Johannesburg, South Africa (an entire line of Witwatersrand gold-mining cities extend east of track [looking right] for hundreds of km), Angolan Biomass Burning (dry-season fires in Angola.  This is the most fire-prone part of the planet due to density of biomass in the equatorial savanna immediately south of the Congo rainforest.  Crew was to look left and right of track), Western Mediterranean Dust (dust should have picked up in North Africa, heading into the western Med ahead of an approaching front), Lower Amazon River Basin (islands and coastal features in the great Amazon estuary should still have been visible between scattered cloud. If visibility permitted, crew was to try for a mapping pass of several overlapping, detailed views), High Central Andean Glaciers (detailed views of any ice-capped volcanoes near nadir. Lenses are now sufficiently powerful that area measurements of snow- and ice-pack can be made from handheld images, using ESC [electronic still camera]), and Lima, Peru (twin city of Lima and its port of Callao just right of track).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 9:35 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (32-amp mode), on backup pump.  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed).  U.S. CDRA CO2 scrubber is on Override.  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.
  • SM Working Compartment:  Pressure (mmHg) — 757, temperature (deg C) — 26.6, ppO2 (mmHg) — 148.6, ppCO2 (mmHg) — 3.6.
  • SM Transfer Compartment:  Pressure (mmHg) — 761, temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 20.7.
  • Node:  Pressure (mmHg) — 749.61, temperature (deg C) — 23.2 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 751.79, temperature (deg C) — 24.9, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 751.79, temperature (deg C) — 25.1; shell heater temp (deg C) — 24.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.3
  • PMA-2:  Shell heater temp (deg C) — 18.3

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  Beta Gimbal Assembly (BGA) 2B and BGA 4B in directed position (Blind mode, non solar-tracking).
  • SM batteries: Battery #1 is off-line, battery #6 is in Cycle mode; all other batteries (6) are in “Partial charge” mode.
  • FGB batteries:  Battery #5 is off-line; all other batteries (5) are in “Partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:

  • C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is operational; PL-2 MDM is Off.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line.
  • State vector — US GPS (SIGI string 1)
  • Attitude — Russian segment
  • Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:

  • All Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-7 viewing position, with Keep Alive power on both strings (based on MBS).
  • MBS:  Keep Alive power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 6:48 am EDT [= epoch]):

  • Mean altitude — 394.7 km
  • Apogee — 406.6 km
  • Perigee — 382.7 km
  • Period — 92.4 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0017647
  • Orbits per 24-hr. day — 15.57
  • Altitude decrease — 150 m (mean) in last 24 hours
  • Solar Beta Angle — -4.3 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 21295
  • Current Flight Attitude — LVLH (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7.8 deg, roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.